If water is to be added to 40 quartz of milk containing 4% of butter fat so as to produce a mixture containing 3% butter fat, How many quartz of water must be added?
Thank you for your help..=)

Answer by stanbon(48535)

If water is to be added to 40 quarts of milk containing 4% of butter fat so as to produce a mixture containing 3% butter fat, How many quarts of water must be added?
---------------------------
Equation:
fat + fat = fat
0.04(40) + 0*x = 0.03(40+x)
Multiply thru by 100 to get:
4(40) = 3(40+x)
160 = 120 + 3x
40 = 3x
x = 40/3 = 13 1/3 quarts of water must be added
=====================================================
Cheers,
Stan H.